# Next Generation LLRF Control Platform for Compact C-band Linear Accelerator

### v1

@@ 1 Introduction
We have investigated RF performance of the RFSoC devices in different frequency ranges and developed readout and control platforms for a number of physics experiments liu2021characterizing ; liu2022development ; liu2023evaluating ; henderson2022advanced ; liu2023higher . For the accelerator applications, the new proposed lepton collider Higgs factory, Cool Cooper Collider (C3), was the initial targeted application for the prototype RFSoC based LLRF Bai:2021rdg . C3 has been proposed as compact and affordable machine and an RFSoC-based LLRF with high integration level can help to achieve the design goals Vernieri:2022fae . Based on the requirements of C3, we have performed a range of targeted configuration optimizations and performance characterizations and summarized the test results in liu2024direct . The RFSoC evaluation board with our custom firmware designs demonstrated pulse to pulse phase jitter as low as 87.54 femtoseconds (fs), which is considerably better than the 150 fs requirement of C3 nanni2023status .
@@ 2 NG-LLRF System and Hardware
Figure 2 shows the hardware setup of the prototype of NG-LLRF. The prototype is developed based on a ZCU216 evaluation board with an UltraScale+ XCZU49DR-2FFVF1760 RFSoC. As Figure 2 shows, the platform has three rf inputs and one rf output. The rf pulse generated by the rf output is amplified by an SSA and then drives the klystron. The high power rf pulse from klystron is coupled to the C3 cavity structure in a test bunker via a waveguide. The resonance frequency of the C3 structure is approximately 5.71209 GHz, which is the rf frequency used for this exercise. The rf signals from the couplers at multiple stages in the test structures are looped back to the rf input channels of the NG-LLRF platform for measurement. For this NG-LLRF experiment, the maximum peak power delivered to the accelerating structure was as high as 16.45 MW. Attenuators are inserted to reduce the power level of signals down to 0 dBm maximum, which is below the 1 dBm maximum input power of ADCs in RFSoC.

### v2

@@ 1 Introduction
We have investigated RF performance of the RFSoC devices in different frequency ranges and developed readout and control platforms for a number of physics experiments liu2021characterizing ; liu2022development ; liu2023evaluating ; henderson2022advanced ; liu2023higher ; liu2024development . For the accelerator applications, the new proposed lepton collider Higgs factory, Cool Cooper Collider (C3), was the initial targeted application for the prototype RFSoC based LLRF Bai:2021rdg . C3 has been proposed as compact and affordable machine and an RFSoC-based LLRF with high integration level can help to achieve the design goals Vernieri:2022fae . Based on the requirements of C3, we have performed a range of targeted configuration optimizations and performance characterizations and summarized the test results in liu2024direct . The RFSoC evaluation board with our custom firmware designs demonstrated pulse to pulse phase jitter as low as 87.54 femtoseconds (fs), which is considerably better than the 150 fs requirement of C3 nanni2023status .
@@ 2 NG-LLRF System and Hardware
Figure 2 shows the hardware setup of the prototype of NG-LLRF. The prototype is developed based on a ZCU216 evaluation board with an UltraScale+ XCZU49DR-2FFVF1760 RFSoC zcu216 . As Figure 2 shows, the platform has three rf inputs and one rf output. The rf pulse generated by the rf output is amplified by an SSA and then drives the klystron. The high power rf pulse from klystron is coupled to the C3 cavity structure in a test bunker via a waveguide. The resonance frequency of the C3 structure is approximately 5.71209 GHz, which is the rf frequency used for this exercise. The rf signals from the couplers at multiple stages in the test structures are looped back to the rf input channels of the NG-LLRF platform for measurement. For this NG-LLRF experiment, the maximum peak power delivered to the accelerating structure was as high as 16.45 MW. Attenuators are inserted to reduce the power level of signals down to 0 dBm maximum, which is below the 1 dBm maximum input power of ADCs in RFSoC.